# Apparently ultra-long period radio sources from self-lensed pulsar-black hole binaries

### v1

@@ 5 The Shapiro delay
Defining a distortion of at least $10\%$ ($\Delta t_{\rm sd}(\Phi_{\rm t})\geq 0.1\,P$) as the significant shape alteration, the threshold orbital phase $\Phi_{\rm th}$ can be derived, within which this distortion criterion holds. The number of pulses that the pulsar can emit when moving from $-\Phi_{\rm th}$ to $\Phi_{\rm th}$ is defined as $n_{\rm sd}=2\Phi_{\rm th}P_{\rm b}/P$. We plot it as a function of $M_{\rm BH}$ in Fig. 6. The blue and red lines correspond to the millisecond and normal pulsar cases, respectively. $n_{\rm lens}=t_{\rm lens}/P$, which is the number of pulses that are affected by the self-lensing effect, is also plotted as dashed lines for comparison. A relatively short orbital period $P_{\rm b}=1\,\rm h$ is chosen, which represents an extreme case compared to most binary systems, and it results in a short $t_{\rm lens}$ and a small $n_{\rm lens}$.
@@ 6 Conclusion
Three recently found long-period radio sources, GLEAM-X J1627 (Hurley-Walker et al., 2022), PSR J0901-4046 (Caleb et al., 2022) and GBM J1839–10 (Hurley-Walker et al., 2023) are puzzling about their physical origins. We interpret them as self-lensed pulsar-BH binaries. Our modeling of them suggests BH masses of $\sim 10^{4}\,\rm M_{\odot}$ and $\sim 4\,\rm M_{\odot}$ for the former two sources, with coalescence times of $\sim 6\times 10^{2}\,\rm yrs$ and $\sim 10^{2}\,\rm yrs$, respectively, basing on their inferred orbital separations. Given the substantial variation of the signal duration, the BH mass of the third source, GBM J1839-10 is estimated to range between $10^{3}$ and $10^{6}\,\rm M_{\odot}$, with a coalescence time spanning $\sim 30$ to $\sim 3\times 10^{3}\,\rm yrs$, implying at least 30 years more of activity, which could be verified by future observation.
[…]
@@
We estimate that, in the Milky Way, there would be at least $80$ NS-BH binaries that can reach a threshold amplification of $10$. Observationally, there would be $\sim 20$ pulsar-BH binaries with self-lensing signals observable for a telescope with a sensitivity of $10\,mJy$, and $\sim 16$ of them would exhibit ultra-long period radio signals. Because of the assumption that all pulsars have a uniform radio luminosity, the actual number of observable sources may be even lower due to the existence of numerous candidates with luminosity lower than the value selected. However, telescopes with lower sensitivities ($\sim\rm\mu Jy$ for long period signals), such as FAST (Han et al., 2021) and SKA (Braun et al., 2019), hold the potential to identify more sources that contain a low luminosity pulsar, increasing the possibility of detecting such systems.
@@
Additionally, we investigate the impact of Shapiro delay on the self-lensed pulse profiles. We find that in pulsar-BH binaries with a normal pulsar, the effect of the Shapiro delay is negligible unless the BH mass exceeds $10^{3}\,\rm M_{\odot}$. However, for systems with a millisecond pulsar, even a stellar-mass BH can alter the shape of the pulse profiles, which would cause at least $10\%$ variation in profile width of the sub-pulses that is emitted inside a threshold orbital phase $\Phi_{\rm th}$, exhibiting either the compression or stretching of the profile.
@@
This work is supported by National Natural Science Foundation of China (grants 12073091, 12261141691 and 12393814). We would like to thank the comments from Baixuan Tan and Long Ji, which have helped improve the quality of the manuscript.

### v2

@@ 5 The Shapiro delay
Defining a distortion of at least $10\%$ ($\Delta t_{\rm sd}(\Phi_{\rm t})\geq 0.1\,P$) as the significant shape alteration, the threshold orbital phase $\Phi_{\rm th}$ can be derived, within which this distortion criterion holds. The number of pulses that the pulsar can emit when moving from $-\Phi_{\rm th}$ to $\Phi_{\rm th}$ is defined as $n_{\rm sd}=2\Phi_{\rm th}P_{\rm b}/P$. We plot it as a function of $M_{\rm BH}$ in Fig. 6. The blue and red lines correspond to the millisecond and normal pulsar cases. $n_{\rm lens}=t_{\rm lens}/P$, which is the number of pulses that are affected by the self-lensing effect, is also plotted as dashed lines for comparison. A relatively short orbital period $P_{\rm b}=1\,\rm h$ is chosen, which represents an extreme case compared to most binary systems, and it results in a short $t_{\rm lens}$ and a small $n_{\rm lens}$.
@@ 6 Discussion and Conclusion
As the self-lensing effect only happens when the binary is edge-on and when the pulsar’s emission cone points toward the observer, the number of non-lensed pulsar-BH binaries should be much more than those self-lensed. However, it usually needs a few years of follow-up observation and precise measurements to determine a non-lensed pulsar-BH binary (Liu et al., 2014). Thus, even if the existing number of the non-lensed systems is large, only a tiny fraction of them can be detected by current observational technique, and it is reasonable that there is no finding of the pulsar-BH binary so far. On the other hand, the self-lensing effect not only makes the signals brighter but also brings them some special observational appearances (see §4), which makes them more easily identified than those non-lensed signals.
@@
We estimate that, in the Milky Way, there would be at least $12$ NS-BH binaries that can reach a threshold amplification of $10$. Observationally, there would be $\sim 3$ pulsar-BH binaries with self-lensing signals observable for a telescope with a sensitivity of $10\,mJy$, and $\sim 2$ of them would exhibit ultra-long period radio signals. Because of the assumption that all pulsars have a uniform radio luminosity, the actual number of observable sources may be even lower due to the existence of numerous candidates with luminosity lower than the value selected. However, telescopes with lower sensitivities ($\sim\rm\mu Jy$ for long period signals), such as FAST (Han et al., 2021) and SKA (Braun et al., 2019), hold the potential to identify more sources that contain a low luminosity pulsar, increasing the possibility of detecting such systems.
Three recently found long-period radio sources, GLEAM-X J1627 (Hurley-Walker et al., 2022), PSR J0901-4046 (Caleb et al., 2022) and GPM J1839-10 (Hurley-Walker et al., 2023) are puzzling about their physical origins. We interpret them as self-lensed pulsar-BH binaries. Our modeling of them suggests BH masses of $\sim 10^{4}\,\rm M_{\odot}$ and $\sim 4\,\rm M_{\odot}$ for the former two sources, with coalescence times of $\sim 6\times 10^{2}\,\rm yrs$ and $\sim 10^{2}\,\rm yrs$, respectively, basing on their inferred orbital separations. Given the substantial variation of the signal duration, the BH mass of the third source, GBM J1839-10 is estimated to range between $10^{3}$ and $10^{6}\,\rm M_{\odot}$, with a coalescence time spanning $\sim 30$ to $\sim 3\times 10^{3}\,\rm yrs$, implying at least 30 years more of activity, which could be verified by future observation.
Our model predicts negative, large values ($\sim 10^{-8}$ s s-1) of the period derivative for the three sources. However, their observed $\dot{P}$’s are all much smaller ($\lesssim 10^{-9}$ s s-1) and even favour positive values (slowing down) for the source PSR J0901-4046. This discrepancy suggests that other physical processes might be involved. For instance, the spin angular momentum of the pulsar might be transferred to the orbital one, via tidal or some sort of magnetic braking effect, thus reversing the orbital period decay.
However, if the three sources are self-lensed NS-BH binaries, their corresponding merger rates derived from their coalescence times would be high ($\sim 10^{4}\,\rm MWEG^{-1}\,Myr^{-1}$), much higher than that given by the up-to-date GW observations ($\sim 5.5\,\rm MWEG^{-1}\,Myr^{-1}$). Therefore, we tend to suggest that the self-lensing scenario for the three sources is disfavoured.
@@
We also investigate the impact of Shapiro delay on the self-lensed pulse profiles. We find that in pulsar-BH binaries with a normal pulsar, the effect of the Shapiro delay is negligible unless the BH mass exceeds $10^{3}\,\rm M_{\odot}$. However, for systems with a millisecond pulsar, even a stellar-mass BH can alter the shape of the pulse profiles, which would cause at least $10\%$ variation in profile width of the sub-pulses that is emitted inside a threshold orbital phase $\Phi_{\rm th}$, exhibiting either the compression or stretching of the profile.
@@
This work is supported by National Natural Science Foundation of China (grants 12073091, 12261141691 and 12393814). We thank the anonymous reviewer whose comments have helped improve the quality of the manuscript. We also thank the comments from Baixuan Tan and Long Ji.